TIME-RESOLVED MW-MW STUDIES OF LOW TEMPERATURE NH_{3} - He COLLISIONS

\underline{D. R. WILLEY}, R. E. TIMLIN JR., I. A. SULAI, J. B. TROMBLEY, Department of Physics, Allegheny College, Meadville, PA 16335.

We report preliminary results of low temperature (10 - 40 K) experimental studies of NH_{3} in collisions with He using time-resolved MW - MW techniques. In these studies, ammonia is collisionally cooled in a low temperature cell filled with a background of helium. A short (< 1μs) pump pulse is used to excite a particular (J, K ) inversion pair while a second weak source probes nearby inversion transitions. At these low temperatures, however, only a few rotational levels are populated, greatly simplifying the interpretation of the time-resolved data. In fact for para - NH_{3} only the (1, 1) and (2, 2) levels will have significant population at 10 - 40 K. Adding in the inversion doublets, para - NH_{3} is effectively a four-level system at these temperatures, making it in principle feasible to extract approximate state-to-state collision rates.